OpenStudy (wes_is_here):

Which ordered pair is the solution to the system of equations? A.(2, −3) B.(−2, 1) C.(5, −2) D.(−6, 5)

OpenStudy (wes_is_here):

x+y=-1 x-3y=11

OpenStudy (wes_is_here):

@J.Maule

OpenStudy (wes_is_here):

@ryry626

OpenStudy (wes_is_here):

@blow_pop2000

OpenStudy (anonymous):

A i think

OpenStudy (wes_is_here):

really

OpenStudy (wes_is_here):

@Brooklyn_Boiii

OpenStudy (nathalie0918):

it is A.

OpenStudy (wes_is_here):

are u sure

OpenStudy (nathalie0918):

yes I am

OpenStudy (wes_is_here):

ok can u help with 4 more

OpenStudy (nathalie0918):

I will try

OpenStudy (wes_is_here):

Which ordered pair is the solution to the system of equations? 2x-y=-5 x+3y=22 A.(1, 7) B.(4, 9) C.(7, 19) D.(8, 21)

OpenStudy (nathalie0918):

a. (1,7)

OpenStudy (wes_is_here):

Which ordered pair is the solution to the system of equations? -5x-3y=12 x-3y=-6 A.(3, −3) B.(−3, 1) C.(6, 4) D.(9, 5)

OpenStudy (nathalie0918):

B. -3,1

OpenStudy (wes_is_here):

Tickets to the zoo cost $14 for adults and $10 for children. A group of 12 people went to the zoo, and the tickets cost $140. The system of equations models this situation, where x is the number of adults and y is the number of children. x+y=12 14x+10y=140 How many adults and how many children were in the group? A.5 adults and 7 children B.6 adults and 6 children C.7 adults and 5 children D.8 adults and 4 children

OpenStudy (nathalie0918):

sorry i stepped out it is A.

OpenStudy (wes_is_here):

are u sure about ur answers

OpenStudy (wes_is_here):

?????

OpenStudy (wes_is_here):

@jordanpipes

OpenStudy (wes_is_here):

@Megatronprime

OpenStudy (megatronprime):

From what I found.....A